OpenStudy (anonymous):

HELP PLSS Let F(x)=5 over X and g(x)=2x^2+5x What two numbers are not in the domain of f o g?

OpenStudy (anonymous):

what do you think f o g means?

OpenStudy (anonymous):

like the x of both equations?

OpenStudy (anonymous):

no. f o g means f[ g(x) ] <-- so what does this also mean?

OpenStudy (anonymous):

f=g(x)?

OpenStudy (anonymous):

no. If you have f(x), what would, say, f(3) means?

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

exactly. So for every x you see, you replace x for 3, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

using the analogy, what do you now think f[ g(x) ] means?

OpenStudy (anonymous):

like x is the number that u will plug in for g and f

OpenStudy (anonymous):

so in this case, f(x) = 5/x. What would you do to f[ g(x) ] ?

OpenStudy (anonymous):

like my answer choices are all 0, and then Ais -5 over 2 B is -2 over 5 C is 2 over 5 and D is 5 over 2 so I'm confused

OpenStudy (anonymous):

???

OpenStudy (anonymous):

i'm trying to help you get the answer. so f(x) = 5/x, f[ g(x) ] would means 5/g(x) right? for every x in f, you replace x with g(x). yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, if you do so, what is the equation would you have?

OpenStudy (anonymous):

umm idk??

OpenStudy (anonymous):

im so confused???

OpenStudy (anonymous):

well you have f(g(x)) = 5/g(x) correct? what is the equation of g(x)?

OpenStudy (anonymous):

its 2x^2+5x

OpenStudy (anonymous):

good. Then what is 5/g(x) look like?

OpenStudy (anonymous):

5 over 2x^2+5x ???

OpenStudy (anonymous):

excellent. f(g(x)) = 5/(2x^2 + 5x) what would you do next?

OpenStudy (anonymous):

would u cross out the 5??

OpenStudy (anonymous):

no you can not. what do you notice you can do with the denominator?

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

can you?

OpenStudy (anonymous):

how would u do that

OpenStudy (anonymous):

you can't. That's nothing can simplify. What kind of function is the denominator? also known as g(x)

OpenStudy (anonymous):

idk what function it is ?

OpenStudy (anonymous):

it's quadratic function yes?

OpenStudy (anonymous):

yes cuz u put a 0 at the end right?

OpenStudy (anonymous):

what 0?

OpenStudy (anonymous):

like equal to zero

OpenStudy (anonymous):

so you're trying to solve x^2 + 5x = 0 ?

OpenStudy (anonymous):

isn't 2x^2+5x=0

OpenStudy (anonymous):

????

OpenStudy (anonymous):

oh you're right. 2x^2 + 5x = 0

OpenStudy (anonymous):

so what would that be ????

OpenStudy (anonymous):

can it be factor?

OpenStudy (anonymous):

yea x(2x+5)

OpenStudy (anonymous):

very good. x (2x + 5) = 0 what would you do next?

OpenStudy (anonymous):

ohhh its 0, -5 over 2

OpenStudy (anonymous):

very good. 0 and -5/2, but why are they not in the domain though?

OpenStudy (anonymous):

because its negative ??

OpenStudy (anonymous):

???

OpenStudy (anonymous):

no what is g(0) and g(-5/2) ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

well, we got x = 2 and x = -5/2 by setting 2x^2 + 5x = 0. so when you plug 2 or -5/2 back in for x, what should you get?

OpenStudy (anonymous):

wait so the answer is 0,-5 over 2 that is not the domain??

OpenStudy (anonymous):

Correct. 0 and -5/2 are the answers. But I want to know why they're not part of the domain

OpenStudy (anonymous):

because if u plug them in it won't give u x?

OpenStudy (anonymous):

plug 0 or -5/2 in what?

OpenStudy (anonymous):

in g of x

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

correct, so what would g(0) and g(-5/2) equal?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

very good. then what is 5/g(0) of 5/g(-5/2) ?

OpenStudy (anonymous):

0, -5 over 2

OpenStudy (anonymous):

???

OpenStudy (anonymous):

????

OpenStudy (anonymous):

g(0) = 0, g(-5/2) = 0 yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

0

OpenStudy (anonymous):

5/0 is 0?

OpenStudy (anonymous):

yea undefiend

OpenStudy (anonymous):

that's right. When x = 0 or -5/2, g(x) would be 0, but you can *not* divide by 0. Therefor, 0 and -5/2 are not part of the domain.

OpenStudy (anonymous):

ohhhh thank u man how can i give u a medal u helped me a lot!!

OpenStudy (anonymous):

just click "Best Response"